Circular No. 7496

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET 2000 S1 B. W. Koehn, Lowell Observatory, reports the discovery by Brian A.

Skiff of a comet on images taken in the course of the LONEOS program. Available CCD observations:

2000 UT α_{2000} δ_{2000} m_{1} Observer
Sept.24.29782 0^{h}31^{m}11.^{s}97 −14^{o}35′32′′.8 14.8 Skiff
24.32105 0 31 10.83 −14 35 26.0  _{”}
24.34425 0 31 09.47 −14 35 19.6  _{”}
24.36749 0 31 08.30 −14 35 12.6  _{”}
25.86483 0 29 50.73 −14 27 35.7 16.0 Tichý
25.86606 0 29 50.61 −14 27 35.9  _{”}
25.86712 0 29 50.49 −14 27 35.5  _{”}

B. A. Skiff (Lowell Observatory). Measurer B. W. Koehn. 0.59-m LONEOS

Schmidt telescope. Fan-shaped tail 2′ long in p.a. ∼ 250^{o}. M. Tichý, M. Kočer, and P. Jelínek (Kleť). 0.57-m f/5.2 reflector. Object

apparently diffuse with coma diameter ∼ 14′′.

2000 DP_{107}

S. J. Ostro, Jet Propulsion Laboratory (JPL); J.-L. Margot and M. C. Nolan, Arecibo Observatory; L. A. M. Benner, R. F. Jurgens, and J. D. Giorgini, JPL, write: “Goldstone (8560-MHz, 3.5-cm) delay-Doppler and Doppler-only radar observations on Sept. 22 and 23 strongly suggest that 2000 DP_{107} is a binary system. The images show separations up to at least 1 km between the components, which have different sizes and rotation states. Additional radar observations are planned at Arecibo. Follow-up optical observations are desirable.”

Ephemeris from orbital elements on MPEC 2000-S37:

2000TT α_{2000} δ_{2000} Δ r ϵ β V
Sept. 26 21^{h}22.^{m}69 −10^{o}41.^{′}3 0.066 1.052 136.^{o}3 41.^{o}2 13.8
27 21 34.85 − 7 45.1 0.072 1.057 138.9 38.6 13.9
28 21 45.26 − 5 12.2 0.077 1.063 140.9 36.4 14.1
29 21 54.25 − 2 59.6 0.083 1.069 142.5 34.8 14.2
30 22 02.09 − 1 04.3 0.089 1.075 143.8 33.4 14.3
Oct. 1 22 08.99 + 0 36.1 0.095 1.080 144.7 32.4 14.4
2 22 15.10 + 2 04.1 0.102 1.086 145.4 31.5 14.6
3 22 20.57 + 3 21.4 0.108 1.092 145.9 30.9 14.7
2000 September 25  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2000 CBAT  Daniel W. E. Green